Calculate the vapor pressure (in torr) at 293 K in a solution prepared by dissolving 19.28 g of the non-volatile non-electrolye urea {CO(NH2)2} in 195 g of water. The vapor pressure of water at 293 K is 17.54 torr. Enter your answer to 2 decimal places.
Answer #1

Given,

The vapour pressure of water at 293 K = 17.54 torr

Mass of urea(non volatile solute) = 19.28g

Mass of the water ( Solvent) = 195g

We have to calculate the vapour pressure of solution at 293 K ,

According to Roult's law,

The vapour pressure of solution of non volatile solute is equal to the vapour pressure of the pure solvent at that temperature multiplied by it's mole fraction.

no of moles= Given mass /Molar mass

Molar mass of urea = 60

So, No of moles of solute (Urea)= 19.28÷60 = 0.32  

Molar mass of water =18

So no of moles of solvent (water) =195÷18 = 10.83

.•. mole fraction of solvent= no of moles of solvent/no of moles of solute+ no of moles of solvent

Mole fraction of solvent= 10.83÷(10.83+0.32) = 10.83÷11.15= 0.97

By applying Roult's law we an can calculate the vapour pressure of solution at 293 K

So , Vapour pressure of solution at 293K= mole fraction of solvent×vapour pressure of pure solvent at 293K

= 0.97 ×17.54= 17.01 torr

Answer , Vapour pressure of solution at 293K is 17.01 torr.